A farmer has 120 meters of fencing and wants to enclose a rectangular plot of land that requires fencing on only three sides since it is bounded by a river on one side. Find the length and width of the plot that will maximize the area?
Can you please help me out? Thanks so much in advance:)

Answer by TimothyLamb(4379)

L + 2w = 120
L = 120 - 2w
---
A = Lw
A = (120 - 2w)w
---
Area is now a quadratic function of width (w):
A(w) = -2ww + 120w + 0
A(w) = (-2)w^2 + 120w + 0
---
the above quadratic equation is in standard form, with a=-2, b=120, and c=0
---
to find the maximum area A(w), plug this:
-2 120 0
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
Answer 1:
the maximum point of the above quadratic equation is: ( 30, 1800 )
---
so the maximum area that can be enclosed is: 1800 sq.m (the y-coordinate of the maximum point)
---
now, to find the L and w dimensions of the maximum area, remember this equation from above:
A(w) = (-2)w^2 + 120w + 0
---
we used that eqn to maximize Area, but as a side effect we also got w for that max area, from the maximum point: ( w=30, A=1800 )
---
Answer 2:
w for maximum area = 30
L for maximum area = 120 - 2w = 120 - 60 = 60
so the maximum area has dims: w=30, L=60
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php